OpenStudy (anonymous):

The per-copy cost of printing, marketing, and shipping x million copies of Danielle Tinne's new novel is given by the formula C(x)=753-225x+3x^3 Find the number of copies that lead to the lower cost per copy, and find that minimum cost.

OpenStudy (radar):

Take the derivative and set to 0, then apply max-min criteria. select appropriate answer.

OpenStudy (anonymous):

please help me

OpenStudy (radar):

C(x) = 3x^3 - 225x + 753 C'(x) = 9x^2 - 225 Set to = 0 in accordance with max/min 9x^2 - 225 = 0 (3x + 15)(3x- 15= 0 Factor in accordance with the difference of two squares.. Solve for x getting x = 5 and x = -5 Only one can be the correct answer, your decision, use logic.